OpenStudy (anonymous):

what equation is equal to 7 times 2/3

OpenStudy (anonymous):

the questions are 5 times 1/3 = 5/3 9 times 1/3 = 9/3 14 times 2/3 = 28/3 14 times 1/3 =14/3

OpenStudy (anonymous):

these are all seperate

OpenStudy (cggurumanjunath):